OpenStudy (anonymous):

the equation of a line is 2y=4x-10. two points that lie on the graph are ?

OpenStudy (anonymous):

just sub in a value for x and solve for y, you can choose whatever x value you want

OpenStudy (anonymous):

what are you saying?

OpenStudy (anonymous):

pick an x value, anyone you want example, 1 or 2 or 3 million pick one

OpenStudy (anonymous):

okay 3

OpenStudy (anonymous):

2y=4x-10 --> 2y=4(3)-10 so 2y=2 so y=??

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so one point would be (3,1) pick another x value